### v1

@@ III-B Constructing Pragmatic Test Set
This test set is designed to challenge LLMs across various conversational scenarios, assessing their ability to interpret implicature in a manner akin to human understanding. By analyzing LLMs’ responses to the test set, we aim to quantify the level of their pragmatic competence, offering insights into both their capabilities and limitations in processing human language. The complete test set is available at https://anonymous.4open.science/r/pragmatic_eval_korean-43D0/.
TABLE III: Experimental Setups for Assessing LLMs’ In-Context Learning Capabilities in the MCQ Test
@@
0-shot (Base) 0 No
1-shot (Base) 1 No
4-shots (Base) 4 No
0-shot (CoT) 0 Yes
1-shot (CoT) 1 Yes
4-shots (CoT) 4 Yes
III-C Multiple-Choice Questions vs. Open-ended Questions
Table II illustrates the variety of question types accompanying each given context and statement within a test unit, categorized by their format into either MCQs or OEQs. MCQs require selecting the most appropriate meaning of the statement among the provided options. Conversely, OEQs involve narratively describing the most appropriate meaning of the statement, without preset options. LLMs are tested in both setups for analyzing differences in performance and behavioral patterns.
For OEQs, we engage three independent assessors who meet the same qualifications as those involved in constructing the test set, to qualitatively assess the narrative responses from LLMs. The assessors rate the LLMs’ responses on a 1 to 5 scale as follows: A score of 5 signifies perfect contextual understanding and accurate interpretation of implicature. A score of 4 reflects an understanding of context with incomplete implicature interpretation. A score of 3 indicates a tendency to focus on the literal meaning, while a score of 2 suggests some misinterpretation of the literal meaning. Lastly, a score of 1 represents a comprehensive misunderstanding of both the context and the literal meaning. For ease of comparison with MCQ outcomes, we present these results as re-scaled scores ranging from 0 to 100. We ensured that informed consent was obtained from all human subjects involved in our study.
III-D In-Context Learning
Recent research highlights the effectiveness of in-context learning, enabling LLMs to rapidly adapt to unfamiliar tasks by leveraging their extensive pre-trained knowledge, without the need for prior training on those tasks [56], [57]. This study focuses on two specific in-context learning strategies: few-shot learning [29] and CoT prompting [30]. Detailed in Table III, our experiment utilizes the MCQ test format to investigate the impact of these strategies on the LLMs’ pragmatic competence, comparing six distinct experimental setups.
We delineate three few-shot learning scenarios: zero-shot, one-shot, and four-shots, each defined by the number of examples provided. The one-shot scenario includes a single test example unit representing the maxim of quality. In the four-shot scenario, four distinct example units are provided, each exemplifying a different maxim, thereby enriching the context provided to LLMs.
Our experiment further distinguishes between two setups based on the implementation of CoT prompting. The base setup omits the CoT approach, presenting solely the test examples and their respective answers. Conversely, the CoT setup includes detailed, step-by-step reasoning processes for the provided examples, encouraging LLMs to explicitly articulate its inferential process as shown in its few-shot examples.
nanummj Specifically, for the zero-shot scenario with CoT, we adopt the methodology of Kojima et al. [58], by appending the phrase “답: 순차적으로 생각해봅시다.” (translated as “Ans- wer: Let’s think step by step.”) at the end of the question, guiding the model towards a structured inferential approach.
IV Experiment
IV-A Experimental Setup
Model Selection
Our study considers five distinct LLMs to compare: GPT-3.5-turbo [59] (henceforth GPT-3.5), GPT-4 [60], Gemini-Pro [61], HyperCLOVA X [48], and LDCC-Solar-10.7B [62] (hereinafter LDCC-Solar). GPT-3.5 and GPT-4 are developed by OpenAI, and Gemini-Pro by Google DeepMind. These models are multilingual LLMs, capable of processing Korean among other languages. In contrast, HyperCLOVA X is developed by NAVER, a leading Korean IT company, and LDCC-Solar is the top-performing open-source model on the Open Ko-LLM Leaderboard as of February 2024, instruction-tuned by Lotte Data Communication using Solar [63] as the foundation model. Both HyperCLOVA X and LDCC-Solar are LLMs specifically optimized for processing Korean.
Hyperparameter Setting
For each LLM, we uniformly configured the hyperparameter settings to ensure a fair comparison across the board. Specifically, the maximum output length was set to 512 tokens, and the temperature parameter was adjusted to 0.7. For HyperCLOVA X, which was not accessible via API, we manually retrieved each response from its website, initiating a new session for each interaction to maintain consistency with the other LLMs.
Performance Report
To report the performance of each LLM, we generated three responses from each model for each test unit. In the case of the MCQ tests, we varied the order of the provided options to mitigate potential bias towards any specific answer position. We then report the averaged scores from three trials to ensure the reliability of our results, taking into account the variability in the outputs of the LLMs.
TABLE IV: Scores on MCQ Test Across four Gricean Maxims
@@ IV-A Experimental Setup
IV-B Result of the MCQ Test
Analysis of LLM Performance
Table IV illustrates the performance of each LLM on the MCQ test across the four Gricean maxims, along with the overall average score. GPT-4 leads with an impressive average score of 81.11, significantly outperforming all compared LLMs. Following closely are HyperCLOVA X and Gemini-Pro, with scores of 67.5 and 59.72, respectively, showcasing their proficiency in understanding conversational implicature.
Notably, LDCC-Solar, despite possessing nearly half the parameter size of GPT-3.5 (10.7 billion compared to GPT-3.5-turbo’s 20 billion) [64], manages to exceed GPT-3.5’s performance by 9.44 points. This highlights the effectiveness of LLMs specialized for Korean, suggesting that a larger parameter size does not necessarily guarantee enhanced performance.
Conversely, the significant performance gap between GPT-3.5 and GPT-4, which boasts 1.7 trillion parameters, underscores the continued importance of parameter scale. This difference serves to emphasize that, while specialized training is crucial, the scale of parameters remains a critical factor in facilitating a model’s capabilities, particularly in tasks requiring nuanced pragmatic inference.
Notably, HyperCLOVA X’s superior performance in the maxim of quality, surpassing even GPT-4, can likely be attributed to its language-specific training. This implies the advantage of developing a model with a comprehensive Korean dataset, which enables it to achieve a deeper understanding of intricate linguistic nuances.
@@ IV-B Result of the MCQ Test
LLMs’ Answer Selection Examined
LDCC-Solar, in particular, shows a tendency to select options not provided in the prompt—such as ‘5’ or ‘d’, which it generates on its own, at a rate of 3.9%—underscoring a significant challenge in adhering to the given choices and further deviating from accurate pragmatic inference. Additionally, LDCC-Solar’s responses often include ‘garbage text’, irrelevant material that detracts from the response quality, pointing to challenges that go beyond understanding pragmatic cues.
IV-C Result of the OEQ Test
TABLE V: Scores on OEQ Test Across four Gricean Maxims
@@ IV-C Result of the OEQ Test
Performance Evaluation from MCQs to OEQs
In Table V showcasing the results of the OEQ test, GPT-4 maintains its leading performance, with HyperCLOVA X and Gemini-Pro closely following, consistent with the MCQ test findings. However, while the performance gap between GPT-4 and HyperCLOVA X was 13.61 in the MCQ test, it has significantly narrowed to 4.13 in the OEQ test, indicating HyperCLOVA X’s enhanced capability in the narrative setup.
@@
Interestingly, LDCC-Solar exhibits a unique behavior by initially generating options akin to those in the MCQ format and then selecting one, even when asked to respond narratively. This behavior likely results from overfitting to the MCQ format, which predominates the benchmarks used in the Open Ko-LLM Leaderboard. Such a strategy, while innovative, may not always align with the expectations for narrative answers and could reflect a limitation in the model’s adaptability to varied question formats. Additionally, as observed in the MCQ test, LDCC-Solar’s responses often include irrelevant text that diminishes the overall quality of its responses.
Analyzing Score Distributions of LLMs
@@
IV-D Results of In-Context Learning
Figure 2: Box Plots of OEQ Score Distributions by each LLM
Table VI showcases the impact of employing two in-context learning techniques in the MCQ setup: few-shot learning and CoT prompting. Reiterating our findings from the MCQ test, GPT-4 continues to outperform all other LLMs across different setups, with HyperCLOVA X and Gemini-Pro following closely. The application of the few-shot learning strategy leads to incremental improvements in performance for most LLMs in both setups with and without CoT prompting.
@@ IV-D Results of In-Context Learning
Considering that the contexts provided in the evaluation were limited to 1-2 sentences, with information crucial for pragmatic inferences not explicitly delineated at the semantic level, the CoT method arguably obstructs the inherent capabilities of LLMs for pragmatic inference. This highlights a limitation of CoT, especially in scenarios where pragmatic cues are subtly embedded below the semantic level and require nuanced interpretation. While CoT has demonstrated potential in enhancing logical thinking and problem-solving in contexts with explicit statements, its effectiveness for pragmatic inference appears contingent on the depth and type of contextual information provided.
TABLE VI: Scores on MCQ Test Under Few-Shot Learning Conditions with and without Chain of Thought prompting
0-shot 1-shot 4-shots
base CoT base CoT base CoT
GPT-3.5 40.00 35.56 44.72 40.56 56.94 47.50
@@
Gemini-Pro 59.72 59.17 63.06 61.94 66.39 62.78
HyperClova X 67.50 66.67 78.06 62.50 84.14 71.94
LDCC-Solar 49.44 52.50 59.44 0.00 0.00 0.00
nanummj Differently from other LLMs, LDCC-Solar exhibits a slight increase in score for the 0-shot CoT setup, indicating some initial advantages of this strategy. However, its performance drastically declines in the 1-shot with CoT and both 4-shot setups, with and without CoT, where it completely fails to generate the required answers, resulting in scores of 0. This failure appears to stem from LDCC-Solar’s tendency to reiterate the given prompt in its responses, which suggests a limitation in its processing capabilities when faced with increased complexity or specificity in the tasks. In most responses, it either repeats the prompts without adding any substantive content or produces irrelevant single words, such as ’철수’ (Cheol-su, a person’s name used in prompts), or meaningless fragments like ’제’ or ’먼’. These responses highlight the model’s difficulty in moving beyond a straightforward single-shot example, possibly reflecting the limits of its capacity to handle nuanced or layered instructions.
TABLE VII: Examples of Korean Culture-Specific Test Units
ID Type of Maxim
@@
27
Quantity
붕어빵 가게 앞을 지나가던 철수가 영희에게
현금이 있는지 물었고 영희는 다음과 같이 말했다.
‘5만원짜리 밖에 없어’.
As Chul-su was passing by a 붕어빵(bung-eo-bbang)
street stall, he asked Yeong-hee if she had any cash.
Yeong-hee replied,
‘I only have a 50,000 won bill’.
83
Relation
철수 집에 놀러 간 영희는 주방에 많은 귤이 쌓여
있는 것을 보고 귤이 왜 이렇게 많은지 물었고
철수는 다음과 같이 말했다.
‘우리 작은 아버지께서 제주도에 사셔’.
When Yeong-hee visited Chul-su’s house, she saw
many tangerines piled up in the kitchen and asked
why there were so many. Chul-su replied,
‘My uncle lives on Jeju Island’.
V Case Study: In-Depth Analysis of LLM Responses to OEQs
V-A A Closer Look at Strengths and Weaknesses
In our analysis, we uncovered distinct characteristics within the responses of LLMs. Most notably, GPT-4, achieving the highest overall score in our evaluations, consistently delivers a singular, unambiguous answer rather than multiple potential interpretations. This decisiveness stands as a substantial strength, streamlining the decision-making process for end-users by eliminating the need to filter through various possible answers. Moreover, GPT-4’s strategy of providing precise responses, and avoiding the use of uncertain expressions like ‘-으로 예상할 수 있다’ (it can be speculated that) or ‘-일 수 있다’ (may be)—expressions that are frequently encountered in GPT-3.5 and occasionally in responses from Gemini and HyperCLOVA X—further contributes to the reliability of its answers.
@@ V-A A Closer Look at Strengths and Weaknesses
HyperCLOVA X demonstrates its strengths by providing detailed explanations of its reasoning process, significantly facilitating a deeper comprehension for users. Conversely, Gemini-Pro reveals a critical limitation by occasionally offering terse responses without elucidation. An example of this is returning a single-word answer, ‘반어법’ (irony), without offering a supplementary interpretation.
V-B Analyzing Error Patterns in LLM Responses
@@
Our examination of LLM responses has identified error patterns that highlight common interpretive challenges. The most prevalent error arises from the LLMs’ struggle to accurately interpret implicatures despite understanding the context, followed by errors due to purely literal interpretations, especially notable in scenarios related to the maxim of manner. This observation contrasts with our findings from the MCQ answer choice distribution analysis (cf. Section 4.2.2), where errors based on literal interpretations were more dominant. This discrepancy implies that the MCQ format, with its predefined options, may have encouraged LLMs to prefer literal meanings, complicating their ability to achieve accurate pragmatic inferences.
We also found that LLMs struggle with generating the correct interpretations when confronted with synonyms or phonetically similar words. A notable case in our test units involved the word ‘사기’ (pronounced as ‘sagi’), which ambiguously means both ‘to buy something’ and ‘a fraud’. Given that ‘사기’ is the nominalized form of ‘사다’ (‘sada’), meaning ‘to buy’, it was expected to be understood in the context of purchasing. Yet, among the LLMs evaluated, only HyperCLOVA X succeeded in accurately interpreting ‘사기’ as related to buying, whereas the others erroneously took it to mean fraud.
@@
Furthermore, we observed errors associated with formatting in the responses. For example, Gemini-Pro occasionally included markdown formatting symbols into its output directly, like **표현** (translated as ‘expression’), with the intention of emphasizing the text. However, these markers appeared literally in the responses, resulting in inaccurately formatted and potentially confusing answers.
V-C Responses to Cultural-specific questions in Korean

### v2

@@ 3.1 Constructing Pragmatic Test Set
This test set is designed to challenge LLMs across diverse conversational scenarios, assessing their ability to interpret implicature in a manner akin to human understanding. The inherent difficulty of each maxim varies, with the maxim of manner often introducing ambiguity that poses additional challenges Hoffmann (2010). We aim to empirically quantify their level of pragmatic competence, offering insights into their capabilities and limitations in processing human language.
3.2 Multiple-Choice Questions vs. Open-ended Questions
Table 3 outlines two types of evaluation questions: MCQs and OEQs. MCQs test a model’s ability to select the most appropriate meaning from provided options, suitable for automated scoring but limited in assessing deeper generative and inferential skills. In contrast, OEQs demand a narrative response, enabling experts to judge the depth and context appropriateness of the answers. This dual strategy evaluates both the basic comprehension and the more complex generative abilities of LLMs.
For OEQs, three independent assessors with qualifications matching those of the test set creators evaluate LLMs’ narrative responses using a Likert scale from 1 to 5. A score of 5 denotes perfect contextual understanding and accurate interpretation of implicature, whereas a score of 1 indicates a complete misunderstanding of both context and literal meaning. Scores are subsequently re-scaled to a 0-100 range for comparison with MCQ outcomes.
3.3 In-Context Learning
Recent research shows that in-context learning allows LLMs to quickly adapt to new tasks using their pre-existing knowledge base, without prior training Dong et al. (2022); Min et al. (2022). This study examines two specific strategies: few-shot learning Brown (2020) and CoT prompting Wei et al. (2022), as detailed in Table 4. We use the MCQ format to compare the impact of these strategies on LLMs’ pragmatic competence across six different setups.
We define three few-shot learning scenarios based on example quantity: zero-shot, one-shot (one example illustrating the maxim of quality), and four-shot (four examples each demonstrating a different maxim). We also compare two CoT prompting setups: the base setup, which only presents test questions and answers, and the CoT setup, which includes detailed reasoning for each question, prompting LLMs to articulate their inferential processes as demonstrated in the few-shot examples.
{CJK}
[…]
@@ 3.3 In-Context Learning
0-shot (Base) 0 X
1-shot (Base) 1 X
4-shots (Base) 4 X
0-shot (CoT) 0 O
1-shot (CoT) 1 O
4-shots (CoT) 4 O
Table 4: Experimental Setups for Assessing LLMs’ In-Context Learning Capabilities in the MCQ Test
4 Experiment
@@
Table 5: Scores on MCQ Test Across four Gricean Maxims
@@
4.1 Experimental Setup
Model Selection.
Our study compares five LLMs: GPT-3.5-turbo and GPT-4 Achiam et al. (2023) by OpenAI, Gemini-Pro Team et al. (2023) by Google DeepMind, HyperCLOVA X Yoo et al. (2024) by NAVER, and LDCC-Solar Kim (2024) by Lotte Data Communication. GPT-3.5, GPT-4, and Gemini-Pro are multilingual, capable of processing Korean among other languages. HyperCLOVA X and LDCC-Solar, however, are optimized specifically for Korean.
Hyperparameter Setting.
For each LLM, we uniformly configured the hyperparameter settings to ensure a fair comparison across the board. The maximum output length was set to 512 tokens, and the temperature parameter was set to 0.7. For HyperCLOVA X, which was not accessible via API, we manually retrieved each response from its website and initiated a new session for each interaction to maintain consistency with the other LLMs.
Performance Report.
We generated three responses from each LLM for each test unit to report their performance. In the case of the MCQ tests, we varied the order of the provided options to mitigate potential bias towards any specific answer position. We report the averaged scores from three trials to ensure the reliability of our results, considering the variability in the LLMs’ outputs.
@@ 4.1 Experimental Setup
Table 6: Scores on OEQ Test Across four Gricean Maxims
Figure 2: Box Plots of OEQ Score Distributions by each LLM
4.2 Result of the MCQ Test
Analysis of LLM Performance.
Table 5 illustrates the performance of each LLM on the MCQ test across the four Gricean maxims, along with the overall average score. GPT-4 leads with an impressive average score of 81.11, significantly outperforming all compared LLMs. HyperCLOVA X and Gemini-Pro follow closely with scores of 67.5 and 59.72, respectively, showcasing their proficiency in understanding conversational implicature.
Notably, LDCC-Solar, with nearly half the parameter size of GPT-3.5-turbo–10.7 billion compared to the reported 20 billion Singh et al. (2023)– manages to exceed GPT-3.5’s performance by 9.44 points. This highlights the effectiveness of LLMs specialized for Korean and suggests that a larger parameter size does not necessarily guarantee better performance.
Conversely, the significant performance gap between GPT-3.5 and GPT-4, which boasts 1.7 trillion parameters, underscores the continued importance of parameter scale. This difference emphasizes that while specialized training is crucial, the scale of parameters remains a critical factor in facilitating a model’s capabilities, particularly in tasks requiring nuanced pragmatic inference.
Notably, HyperCLOVA X’s superior performance in the maxim of quality, surpassing even GPT-4, can likely be attributed to its language-specific training. This implies the advantage of developing a model with a comprehensive dataset in Korean, which enables it to achieve a deeper understanding of intricate linguistic nuances.
LLMs’ Answer Selection Examined.
LDCC-Solar, in particular, shows a tendency to select options not provided in the prompt—such as ‘5’ or ‘d’, which it generates on its own, at a rate of 3.9%—underscoring a significant challenge in adhering to the given choices and further deviating from accurate pragmatic inference. Additionally, LDCC-Solar’s responses often include noisy text, irrelevant material that detracts from the response quality, pointing to challenges that go beyond understanding pragmatic cues.
4.3 Result of the OEQ Test
Performance Evaluation from MCQs to OEQs.
In Table 6 showcasing the results of the OEQ test, GPT-4 maintains its leading performance, with HyperCLOVA X and Gemini-Pro closely following, consistent with the MCQ test findings. However, while the performance gap between GPT-4 and HyperCLOVA X was 13.61 in the MCQ test, it has significantly narrowed to 4.13 in the OEQ test, indicating HyperCLOVA X’s enhanced capability in the narrative setup.
@@ 4.3 Result of the OEQ Test
Interestingly, LDCC-Solar exhibits a unique behavior by initially generating options akin to those in the MCQ format and then selecting one, even when asked to respond narratively. This behavior likely results from overfitting to the MCQ format, which predominates the benchmarks used in the Open Ko-LLM Leaderboard. Such a strategy, while innovative, may not always align with the expectations for narrative answers and could reflect a limitation in the model’s adaptability to varied question formats. Additionally, as observed in the MCQ test, LDCC-Solar’s responses often include noisy text that diminishes the overall quality of its responses.
0-shot 1-shot 4-shots
base CoT base CoT base CoT
GPT-3.5 40.00 35.56 44.72 40.56 56.94 47.50
GPT-4 81.11 77.22 86.11 80.56 89.17 84.44
Gemini-Pro 59.72 59.17 63.06 61.94 66.39 62.78
HyperClova X 67.50 66.67 78.06 62.50 84.14 71.94
LDCC-Solar 49.44 52.50 59.44 0.00 0.00 0.00
Table 7: Scores on MCQ Test Under Few-Shot Learning Conditions with and without Chain of Thought prompting
Analyzing Score Distributions of LLMs.
@@
4.4 Results of In-Context Learning
Table 7 showcases the impact of employing two in-context learning techniques in the MCQ setup: few-shot learning and CoT prompting. Consistent with our findings from the MCQ test, GPT-4 continues to outperform all other LLMs across different setups, with HyperCLOVA X and Gemini-Pro following closely. The application of the few-shot learning strategy leads to incremental improvements in performance for most LLMs in both setups with and without CoT prompting.
@@ 4.4 Results of In-Context Learning
The context units provided in the evaluation data set were limited to 1-2 sentences, with information crucial for pragmatic inferences not explicitly delineated at the semantic level. In such cases, the CoT method arguably obstructs the inherent capabilities of LLMs for pragmatic inference. This highlights a limitation of CoT, especially in scenarios where pragmatic cues are subtly embedded below the semantic level and require nuanced interpretation. While CoT has demonstrated potential in enhancing logical thinking and problem-solving in contexts with explicit statements, its effectiveness for pragmatic inference appears contingent on the depth and type of contextual information provided.
{CJK}
@@
Id Maxim
@@
27 Quantity 붕어빵 가게 앞을 지나가던 철수가 영희에게 현금이 있는지 물었고 영희는 다음과 같이 말했다. ‘5만원짜리 밖에 없어’. As Chul-su was passing by a 붕어빵(bung-eo-bbang) street stall, he asked Yeong-hee if she had any cash. Yeong-hee replied, ‘I only have a 50,000 won bill’.
83 Relation 철수 집에 놀러 간 영희는 주방에 많은 귤이 쌓여 있는 것을 보고 귤이 왜 이렇게 많은지 물었고 철수는 다음과 같이 말했다. ‘우리 작은 아버지께서 제주도에 사셔’. When Yeong-hee visited Chul-su’s house, she saw many tangerines piled up in the kitchen and asked why there were so many. Chul-su replied, ‘My uncle lives on Jeju Island’.
Table 8: Examples of Korean Culture-Specific Test Units
5 Case Study: In-Depth Analysis of LLM Responses to OEQs
5.1 A Closer Look at Strengths and Weaknesses
GPT-4, the top performer in our evaluations, excels by offering clear, definitive answers rather than multiple interpretations. This simplifies decision-making for users by eliminating the need to filter through various possible answers. Moreover, its avoidance of uncertain expressions such as ‘-으로 예상할 수 있다’ (it can be speculated that) or ‘-일 수 있다’ (may be)—common in GPT-3.5 and occasionally seen in Gemini and HyperCLOVA X—further contributes to the reliability of its answers.
@@ 5.1 A Closer Look at Strengths and Weaknesses
HyperCLOVA X demonstrates its strengths by providing detailed explanations of its reasoning process, significantly facilitating a deeper comprehension for users. Conversely, Gemini-Pro reveals a critical limitation by occasionally offering brief responses without explanation. An example of this is returning a single-word answer, ‘반어법’ (irony), without offering a supplementary interpretation.
5.2 Analyzing Error Patterns in LLM Responses
@@
The most prevalent error arises from struggling to interpret implicatures, despite understanding the context, followed by errors due to purely literal interpretations, especially concerning the maxim of manner. This contrasts with our MCQ answer choice analysis (cf. Section 4.2), where literal interpretation errors were more prevalent. This indicates that the MCQ format may have prompted LLMs to favor literal meanings, hindering accurate pragmatic inferences.
LLMs also struggle with interpreting synonyms or phonetically similar words. A notable example is the word ‘사기’ (‘sagi’), meaning both ‘to buy’ and ‘fraud.’ Despite the context indicating the purchasing meaning, only HyperCLOVA X correctly identified it, while the others misinterpreted it as fraud.
@@
We also observed errors associated with formatting in the responses. For example, Gemini-Pro occasionally included markdown formatting symbols into its output directly, like **표현** (‘expression’), with the intention of emphasizing the text. However, these markers appeared verbatim in the responses, resulting in inaccurately formatted and potentially confusing answers.
5.3 Responses to Cultural-specific questions in Korean